# HSR-Enhanced Sparse Attention Acceleration

### v1

@@ Appendix F Error Analysis of Softmax Attention
Lemma F.1 ( General error analysis of Softmax attention with index set, formal version of Lemma 6.4 ).
@@
  • Let $\alpha,\overline{\alpha}$ and $\widehat{\mathsf{Attn}}_{s}$ be defined as Definition 3.2.
@@
Recall that $\overline{R}=[n]\setminus R$ and $\widehat{K}=K_{R}\in\mathbb{R}^{r\times d}$ and $\widehat{V}=V_{R}\in\mathbb{R}^{r\times d}$ and $\overline{K}=K_{\overline{R}}\in\mathbb{R}^{(n-r)\times d}$ and $\overline{V}=V_{\overline{R}}\in\mathbb{R}^{(n-r)\times d}$ as defined in Definition 3.1. Also, we have $\widehat{u}=\exp(q\widehat{K}^{\top})\in\mathbb{R}^{r}$ and $\widehat{\alpha}=\langle\widehat{u},{\bf 1}_{r}\rangle\in\mathbb{R}$ and $\overline{u}=\exp(q\overline{K}^{\top})\in\mathbb{R}^{n-r}$ and $\overline{\alpha}=\langle\overline{u},{\bf 1}_{n-r}\rangle\in\mathbb{R}$ as defined in Definition 3.2.
@@
where the first step is by Definition 3.2, the second step is by triangle inequality, the third step is by $\|uV\|_{\infty}\leq\|u\|_{1}\cdot\|V\|_{\infty}$ for any vector $u$ and conformable matrix $V$, and the fourth step is by definition of $\widehat{\alpha}$ and $\overline{\alpha}$, i.e., $\widehat{\alpha}=\langle\widehat{u},{\bf 1}_{r}\rangle=\|\widehat{u}\|_{1}$ (note that each entry of $\widehat{u}$ is positive), the fifth step is by $\max\{\|\widehat{V}\|_{\infty},\|\overline{V}\|_{\infty}\}=\|V\|_{\infty}$, the sixth step in by simple calculation and the last step is by $\widehat{\alpha}+\overline{\alpha}=\alpha$. ∎
@@
Theorem F.2 (Error analysis of Softmax attention with index set, formal version of Theorem 4.3).
@@
  • Let the Softmax attention with index set $\widehat{\mathsf{Attn}}_{s}$ be defined as Definition 3.2.
@@
Let $\alpha,\overline{\alpha},\widehat{\alpha}$ be defined in Definition 3.2. By Lemma F.1, we have
@@
where the first step is by Definition of $\widehat{\alpha}$, the second step is by Definition 3.3 and Jensen inequality, and the last step is by simple calculation.
@@
where the first step is by Definition of $\overline{\alpha}$, the second step is by Definition 3.3, and the last step is by simple calculation.

### v2

@@ Appendix G Error Analysis of Softmax Attention
Lemma G.1 ( General error analysis of Softmax attention with index set, formal version of Lemma 6.5 ).
@@
  • Let $\alpha,\overline{\alpha}$ and $\widehat{\mathsf{Attn}}_{s}$ be defined as Definition B.2.
@@
Recall that $\overline{R}=[n]\setminus R$ and $\widehat{K}=K_{R}\in\mathbb{R}^{r\times d}$ and $\widehat{V}=V_{R}\in\mathbb{R}^{r\times d}$ and $\overline{K}=K_{\overline{R}}\in\mathbb{R}^{(n-r)\times d}$ and $\overline{V}=V_{\overline{R}}\in\mathbb{R}^{(n-r)\times d}$ as defined in Definition B.1. Also, we have $\widehat{u}=\exp(q\widehat{K}^{\top})\in\mathbb{R}^{r}$ and $\widehat{\alpha}=\langle\widehat{u},{\bf 1}_{r}\rangle\in\mathbb{R}$ and $\overline{u}=\exp(q\overline{K}^{\top})\in\mathbb{R}^{n-r}$ and $\overline{\alpha}=\langle\overline{u},{\bf 1}_{n-r}\rangle\in\mathbb{R}$ as defined in Definition B.2.
@@
where the first step is by Definition B.2, the second step is by triangle inequality, the third step is by $\|uV\|_{\infty}\leq\|u\|_{1}\cdot\|V\|_{\infty}$ for any vector $u$ and conformable matrix $V$, and the fourth step is by definition of $\widehat{\alpha}$ and $\overline{\alpha}$, i.e., $\widehat{\alpha}=\langle\widehat{u},{\bf 1}_{r}\rangle=\|\widehat{u}\|_{1}$ (note that each entry of $\widehat{u}$ is positive), the fifth step is by $\max\{\|\widehat{V}\|_{\infty},\|\overline{V}\|_{\infty}\}=\|V\|_{\infty}$, the sixth step in by simple calculation and the last step is by $\widehat{\alpha}+\overline{\alpha}=\alpha$. ∎
@@
Theorem G.2 (Error analysis of Softmax attention with index set, formal version of Theorem 4.3).
@@
  • Let the Softmax attention with index set $\widehat{\mathsf{Attn}}_{s}$ be defined as Definition B.2.
@@
Let $\alpha,\overline{\alpha},\widehat{\alpha}$ be defined in Definition B.2. By Lemma G.1, we have
@@
where the first step is by Definition of $\widehat{\alpha}$, the second step is by Definition B.3 and Jensen inequality, and the last step is by simple calculation.
@@
where the first step is by Definition of $\overline{\alpha}$, the second step is by Definition B.3, and the last step is by simple calculation.